# Valid Stopping for LLM Generation via Empirical Dynamic Formal Lift

### v1

@@ Abstract
We introduce Sequential-EDFL (Empirical Dynamic Formal Lift), applying anytime-valid sequential testing to language model generation stopping. Our approach tracks information lift—the log-likelihood ratio between full models and deliberately weakened “skeleton” baselines—using self-normalized empirical-Bernstein e-processes that provide formal $\delta$-level error control regardless of stopping time. We handle unknown centering through online mean estimation, combine multiple parameters via mixture e-processes, and support adaptive resets under distributional drift. On six benchmarks, Sequential-EDFL reduces generation by 22–28% vs. sequential baselines while maintaining $\delta$-level control with 12% computational overhead. We introduce automated skeletons (distilled submodels, randomized logits) and show robustness across skeleton families. Composing EDFL with a lightweight correctness gate (sentence boundaries + verifier) improves end-task correctness while preserving anytime-valid guarantees by only delaying stopping. Our certificates control information sufficiency, not factual correctness—10.9% of stopped sequences remain incorrect even with the gate (13.2–22.7% without it). EDFL serves as a first-stage filter reducing verification burden by 83%, not as a standalone solution for safety-critical domains. [^1] Computed as: gate triggers on 83% of stops per Table 17, requiring verification only on remaining 17% of stops plus all non-stopped sequences.
Valid Stopping for LLM Generation via Empirical Dynamic Formal Lift
Sanjeda Akterthanks: Equal contribution.1  and Ibne Farabi Shihab11footnotemark: 1thanks: Corresponding author: ishihab@iastate.edu.1  and Anuj Sharma2 1Department of Computer Science, Iowa State University 2Department of Civil, Construction & Environmental Engineering, Iowa State University
@@ 1 Introduction
Language models generate sequences token-by-token, yet determining when to stop generation remains an open challenge. Current approaches rely on fixed-length generation, simple heuristics (entropy thresholds, EOS tokens), or ad-hoc stopping criteria without statistical guarantees. Language models often generate extraneous content beyond minimal sufficient answers Xu et al. (2024), while existing stopping methods Wei et al. (2022); Manakul et al. (2023) provide no formal error control. Unlike classical sequential testing with independent samples, language generation exhibits complex dependencies, unknown distributions, and time-varying statistics. We ask: Can we provide anytime-valid statistical guarantees for language model stopping decisions? Recent advances in sequential testing through e-processes Vovk and Wang (2021b); Howard et al. (2021b)—nonnegative martingales that enable hypothesis testing with anytime-valid error control—offer a promising foundation, generalizing likelihood ratios to sequential settings. However, applying e-processes to language generation faces fundamental challenges: unknown conditional expectations, distributional drift, and defining meaningful null hypotheses for “information sufficiency.”
We introduce EDFL with four contributions: (i) a skeleton-based information lift that quantifies evidence accumulation; (ii) self-normalized empirical-Bernstein e-processes and mixture e-processes that yield anytime-valid $\delta$-control without oracle centering; (iii) adaptive segment budgeting to accommodate drift; and (iv) practical add-ons that improve deployability: a verifier-gated stopping variant that only delays stopping (thus preserving anytime validity) and a skeleton diagnostics checklist with task-agnostic, automated skeletons. Across six benchmarks (GSM8K, HotpotQA, ASQA, TruthfulQA, ProofWriter, LegalBench), Sequential-EDFL reduces generation by 22-28% vs. sequential baselines while maintaining $\delta$-level error control with 12% computational overhead. Critically, our certificates guarantee information sufficiency relative to the skeleton baseline, not factual correctness—13.2–22.7% of high-lift sequences produce incorrect answers (Table 9), representing an important limitation for practical deployment. To mitigate skeleton dependence, we propose task-agnostic automated skeletons (distilled or subsampled submodels) and empirically demonstrate robustness to skeleton choice. To narrow the sufficiency–correctness gap, we add a hybrid correctness gate that enforces sentence-boundary stops and triggers a lightweight verifier; this preserves the statistical stopping guarantees while improving factuality in practice.
EDFL certifies information sufficiency with respect to a fixed skeleton; it does not certify factual correctness. Skeleton selection is required; we make it practical via automated recipes (distilled submodels, randomized flattening) plus diagnostics (KL $\in$ [2,10] nats, $\rho<-0.5$). The hybrid gate mitigates but does not eliminate the correctness gap. For safety-critical applications, EDFL serves as a first-stage filter reducing verification load, not as a replacement for domain-critical verification.
@@ 2 Information Lift Framework
Our approach measures information accumulation through skeleton-based lift, comparing full model predictions against deliberately weakened baselines. This section formalizes the lift computation and describes principled skeleton construction methods.
2.1 Skeleton-Based Information Lift
We define information lift as the log-likelihood ratio between full model predictions and skeleton baselines:
@@ 2.1 Skeleton-Based Information Lift
2.2 Skeleton Construction and Diagnostics
Effective skeleton construction requires creating information-poor baselines while maintaining computational efficiency. We employ three principled approaches: (1) Prompt Compression removes contextual information (examples, retrieved documents) while preserving task structure; (2) Context Ablation directly measures information gain from external sources (optimal for RAG); (3) Temperature Scaling ($\tau\in[1.5,2.0]$) flattens output distributions as a general-purpose approach.
We validate skeleton quality through a diagnostics checklist requiring: (i) KL divergence $\mathrm{KL}(P\|S)\in[2,10]$ nats, (ii) negative entropy correlation $\rho(X_{t},H_{t})<-0.5$, and (iii) low clipping saturation (<5<5% tokens). Full construction recipes, validation metrics across all datasets (Table 7), automated skeleton families, and the lift illustration appear in Appendix A.
For deployments requiring task-agnostic baselines, we provide two automated skeleton families: (1) distilled/subsampled submodels and (2) randomized logit flattening. These require no domain expertise and preserve EDFL’s guarantees since $S$ is chosen before testing. Table 8 demonstrates stable performance across skeleton families. Full recipes and validation appear in Appendix B.
@@ 3 Sequential E-Process Framework
We assume bounded, nonnegative increments $X_{t}\in[0,c]$, predictable clipping, and that skeleton $S$ is fixed prior to testing. We briefly roadmap the theory: we (i) define a self-normalized e-process that remains a supermartingale under online mean/variance estimation, (ii) form a mixture e-process over a grid of $\lambda$ values, and (iii) allocate error budgets across segments to handle drift while maintaining global $\delta$-control.
Given skeleton-based information lift $X_{t}$ from Section 2, we construct anytime-valid stopping rules using self-normalized empirical-Bernstein e-processes. This section develops the statistical framework enabling formal $\delta$-level error control regardless of stopping time.
The fundamental challenge lies in unknown centering: we observe $X_{t}$ but not the conditional expectation $\mu_{t}=\mathbb{E}[X_{t}|\mathcal{F}_{t-1}]$. Classical e-processes assume known centering, but language generation exhibits unknown and time-varying statistics. We address this through self-normalized empirical-Bernstein e-processes that estimate conditional means online while maintaining the supermartingale property.
@@
where $\hat{\mu}_{s}$ and $\hat{v}_{s}$ are running estimates of mean and variance. The key insight is that conservative estimation—systematically overestimating variance and underestimating favorable deviations—preserves the supermartingale property even under estimation error.
@@
This theorem provides the formal foundation for our method: we stop generation when accumulated information lift exceeds a statistical boundary, with Type I error controlled at level $\delta$ regardless of when we choose to stop. The proof combines techniques from self-normalized processes Howard et al. (2021a), mixture e-processes Vovk and Wang (2021b), and adaptive budgeting via convergent series.
@@
Proof. See Appendix J.
The resulting boundary $u_{J}$ grows quadratically as $J^{2}$, naturally limiting resets through increasingly stringent stopping thresholds. For instance, with $\delta=0.1$, we have $u_{1}=\pi^{2}/(6\delta)\approx 16.4$ for the first segment and $u_{5}=25\pi^{2}/(6\delta)\approx 411$ for the fifth segment—a 25-fold increase that discourages excessive resets while maintaining statistical validity.
At each token, compute lift $X_{t}$, optionally skip on entropy slope, update $(\hat{\mu}_{t},\hat{v}_{t})$ and $\lambda e-pr$ocesses, mix $M_{t}=\sum_{k}w_{k}M_{t}(\lambda_{k})$, stop when $M_{t}\geq u_{J}$ (with a gate that only delays), and reset on drift with a new budget $\delta_{J}$. Full pseudocode: Appendix N, Algorithm 1.
The optional gate enforces sentence-boundary stops and a lightweight verifier (e.g., retrieval overlap, arithmetic checker, or self-consistency score). Because the e-process threshold $u_{J}$ is unchanged and the gate enforces only monotone delays ($\tau^{\prime}\geq\tau$), the anytime-valid $\delta$-control for sufficiency is preserved (Lemma 3.2). The gate cannot cause earlier stopping, thus cannot increase Type I error. Complete pseudocode and method overview diagram appear in Algorithm 1 (Appendix N).
The VerifierPass function in Algorithm 1 (line 23) is task-specific: For GSM8K, we use symbolic calculator execution on extracted numerical expressions. For HotpotQA/ASQA, we compute retrieval overlap score $\geq 0.3$ (answer spans supported by retrieved documents). For TruthfulQA, we use knowledge base lookup with exact match. For ProofWriter, we apply theorem prover validation. For LegalBench, we check legal database precedent match. The threshold $\tau_{c}$ is set to 0.7 across all tasks based on validation set calibration.
Having established the theoretical framework and algorithmic implementation of Sequential-EDFL, we now empirically evaluate its performance across diverse language generation tasks to validate both the formal guarantees and practical efficiency.
@@ 4 Experiments
We evaluate Sequential-EDFL on generation stopping across six benchmarks: GSM8K Cobbe et al. (2021) (mathematical reasoning), HotpotQA Yang et al. (2018) (multi-hop QA), ASQA Stelmakh et al. (2022) (long-form QA), TruthfulQA Lin et al. (2022) (factual accuracy), ProofWriter (logical reasoning), and LegalBench (legal analysis). We compare against three verification strategies: Verify-All (uniform verification), Random-50% (random sampling), and Confidence-Based (verify when softmax confidence <0.9<0.9).
Models: We evaluate on GPT-4 (gpt-4-0613), LLaMA-2-7B, and LLaMA-2-13B. Verification mechanisms vary by domain: GSM8K uses symbolic calculator execution, HotpotQA and ASQA use search engine validation (top-5 results), TruthfulQA uses knowledge base fact-checking, ProofWriter uses theorem prover validation, and LegalBench uses legal database lookup.
Calibration: For each benchmark, we randomly split examples 70% train, 10% calibration, 20% test. We train the correctness predictor on training data, compute conformal thresholds on calibration data, and report all metrics on held-out test sets. We set $\delta\in\{0.03,0.05,0.10\}$ to examine error-cost tradeoffs.
Metrics: We report: (1) Verification Rate: percentage of outputs verified, (2) Accuracy: correctness rate on all outputs (verified outputs use verified answers, skipped outputs use model answers), (3) Cost Reduction: $(1-\text{VR})\times 100\%$ where VR is verification rate, (4) False Negative Rate: percentage of incorrect outputs that skipped verification. We report TPCA (Tokens Per Certified Answer) as our primary efficiency metric; TPCA is defined as the mean number of generated tokens until a certification event (lower is better).
We implement the standard e-value stopping rule using a fixed-$\lambda$ non-mixture e-process (no self-normalization), following the betting/e-process literature (Vovk and Wang, 2021b; Ramdas et al., 2023). We sweep $\lambda$ on a validation split and report test performance.
4.1 Main Results: Token Reduction with Formal Guarantees
Table 1 presents our primary results comparing stopping methods across benchmarks. Sequential-EDFL achieves substantial token reduction while maintaining $\delta$-level error control with moderate computational overhead.

### v2

# Anytime-Valid Answer Sufficiency Certificates for LLM Generation via Sequential Information Lift

@@ Abstract
We introduce Sequential-EDFL (Empirical Dynamic Formal Lift), applying anytime-valid sequential testing to language model generation stopping. Our approach tracks information lift (the log-likelihood ratio between full models and deliberately weakened “skeleton” baselines) using self-normalized empirical-Bernstein e-processes that provide formal $\delta$-level error control regardless of stopping time. The $\delta$ guarantee controls premature stopping with insufficient information lift relative to the skeleton, and does not imply $\delta$ control of factual incorrectness or hallucinationsakter2025inducing. We handle unknown centering through online mean estimation, combine multiple parameters via mixture e-processes, and support adaptive resets under distributional drift. On six benchmarks, Sequential-EDFL reduces generation by 22–28% vs. sequential baselines while maintaining $\delta$-level control with 12% computational overhead. We introduce automated skeletons (distilled submodels, randomized logits) and show robustness across skeleton families. Composing EDFL with a lightweight correctness gate (sentence boundaries + verifier) improves end-task correctness while preserving anytime-valid guarantees by only delaying stopping. Our certificates control information sufficiency, not factual correctness. Specifically, 10.9% of stopped sequences remain incorrect even with the gate (13.2–22.7% without it). EDFL serves as a first-stage filter that can reduce verification burden (when applied to stopped sequences, the gate validates 83% of stops, requiring full verification only for the remaining 17% plus all non-stopped sequences), not as a standalone solution for safety-critical domains. [^1] Computed as: gate triggers on 83% of stops per Table 3, requiring verification only on remaining 17% of stops plus all non-stopped sequences. The overall reduction depends on the stopping rate $p_{\text{stop}}$: reduction = $0.83\cdot p_{\text{stop}}$.
Anytime-Valid Answer Sufficiency Certificates for LLM Generation via Sequential Information Lift
Ibne Farabi Shihabthanks: Equal contribution.thanks: Corresponding author: ishihab@iastate.edu.1  and Sanjeda Akter11footnotemark: 11  and Anuj Sharma2 1Department of Computer Science, Iowa State University 2Department of Civil, Construction & Environmental Engineering, Iowa State University ishihab@iastate.edu
@@ 1 Introduction
Determining when to stop language model generation remains an open challenge. Current approaches rely on fixed-length generation or simple heuristics without statistical guarantees wei2022chain; manakul2023selfcheckgpt, while models often generate extraneous content xu2024hallucination. Unlike classical sequential testing, language generation exhibits complex dependencies and unknown distributions. We ask: Can we provide anytime-valid control of premature stopping? E-processes vovk2021evalue; howard2021time, nonnegative martingales enabling anytime-valid hypothesis testing, offer a promising foundation but require handling unknown conditional expectations and distributional drift.
We introduce Sequential-EDFL (Empirical Dynamic Formal Lift) with four contributions. We provide anytime-valid $\delta$-control for information sufficiency via skeleton-based information lift and self-normalized empirical-Bernstein e-processes, guaranteeing that with probability at least $1-\delta$ we do not stop prematurely before sufficient information accumulation (Theorem 3.1). This is not a guarantee of factual correctness. We develop adaptive segment budgeting with convergent-series error allocation to handle distributional drift. We provide principled skeleton construction with automated recipes and diagnostics requiring no domain expertise. We demonstrate practical systems integration via a hybrid correctness gate that preserves anytime-valid guarantees while improving factuality.
On six benchmarks, Sequential-EDFL reduces generation by 22-28% vs. heuristic baselines while maintaining $\delta$-level sufficiency control with 12% overhead. Our certificates guarantee information sufficiency relative to skeleton baselines, not factual correctness: 13.2–22.7% of high-lift sequences produce incorrect answers (Table 10). The hybrid gate reduces this gap to 10.9% but remains insufficient for safety-critical applications without complementary verification. For such applications, EDFL serves as a first-stage filter, not a replacement for domain-critical verification.
@@ 2 Information Lift Framework
Our approach measures information accumulation through skeleton-based lift, comparing full model predictions against deliberately weakened baselines. We define information lift as the log-likelihood ratio between full model predictions and skeleton baselines:
@@
We validate skeleton quality through diagnostics (KL divergence $\in$ [2,10] nats, negative entropy correlation) and provide automated skeleton families for task-agnostic deployment (Appendices B and C).
@@ 3 Sequential E-Process Framework
Given skeleton-based information lift $X_{t}$ from Section 2.1, we construct anytime-valid stopping rules using self-normalized empirical-Bernstein e-processes. We assume bounded, nonnegative increments $X_{t}\in[0,c]$, predictable clipping, and that skeleton $S$ is fixed prior to testing. We briefly roadmap the theory: we define a self-normalized e-process that remains a supermartingale under online mean/variance estimation, form a mixture e-process over a grid of $\lambda$ values, and allocate error budgets across segments to handle drift while maintaining global $\delta$-control.
The fundamental challenge lies in unknown centering: we observe $X_{t}$ but not the conditional expectation $\mu_{t}=\mathbb{E}[X_{t}|\mathcal{F}_{t-1}]$. Classical e-processes assume known centering, but language generation exhibits unknown and time-varying statistics akter2025selective. We address this through self-normalized empirical-Bernstein e-processes that estimate conditional means online while maintaining the supermartingale property.
@@
where $\hat{\mu}_{s}$ and $\hat{v}_{s}$ are running estimates of mean and variance computed via exponential moving averages (EMAs) with rate $\alpha$. The estimates must be predictable (measurable w.r.t. $\mathcal{F}_{s-1}$), which we ensure by using $\hat{\mu}_{s-1}$ and $\hat{v}_{s-1}$ in the exponent at step $s$. The key insight is that conservative estimation, which systematically overestimates variance (via inflation factor $\eta>0$) and uses EMA-based smoothing, preserves the supermartingale property under bounded estimation error. We validate this conservatism empirically (Section 4).
@@
This theorem provides the formal foundation for our method: we stop generation when accumulated information lift exceeds a statistical boundary, with Type I error (premature stopping with insufficient lift) controlled at level $\delta$ regardless of when we choose to stop. The proof combines techniques from self-normalized processes howard2021timeuniform, mixture e-processes vovk2021evalue, and adaptive budgeting via convergent series. Operational interpretation: The guarantee means that with probability at least $1-\delta$, we do not stop before information lift has accumulated sufficiently relative to the skeleton baseline. This does not guarantee factual correctness (see Section 4.2).
@@
Proof. See Appendix L.
The resulting boundary $u_{J}$ grows quadratically as $J^{2}$, naturally limiting resets through increasingly stringent stopping thresholds. For instance, with $\delta=0.1$, we have $u_{1}=\pi^{2}/(6\delta)\approx 16.4$ for the first segment and $u_{5}=25\pi^{2}/(6\delta)\approx 411$ for the fifth segment, a 25-fold increase that discourages excessive resets while maintaining statistical validity.
At each token, compute lift $X_{t}$, optionally skip on entropy slope, update $(\hat{\mu}_{t},\hat{v}_{t})$ and $\lambda e-pr$ocesses, mix $M_{t}=\sum_{k}w_{k}M_{t}(\lambda_{k})$, stop when $M_{t}\geq u_{J}$ (with a gate that only delays), and reset on drift with a new budget $\delta_{J}$. Full pseudocode: Appendix O, Algorithm 1.
The optional gate enforces sentence-boundary stops and a lightweight verifier (e.g., retrieval overlap, arithmetic checker, or self-consistency score). Because the e-process threshold $u_{J}$ is unchanged and the gate enforces only monotone delays ($\tau^{\prime}\geq\tau$), the anytime-valid $\delta$-control for sufficiency is preserved (Lemma 3.2). The gate cannot cause earlier stopping, thus cannot increase Type I error. Complete pseudocode and method overview diagram appear in Algorithm 1 (Appendix O).
The VerifierPass function (Algorithm 1, line 23) is task-specific: symbolic calculator (GSM8K), retrieval overlap $\geq 0.3$ (HotpotQA/ASQA), knowledge base lookup (TruthfulQA), theorem prover (ProofWriter), legal database match (LegalBench), with threshold $\tau_{c}$ = 0.7 validated on calibration sets. Having established the theoretical framework and algorithmic implementation of Sequential-EDFL, we now empirically evaluate its performance across diverse language generation tasks to validate both the formal guarantees and practical efficiency.
@@ 4 Experiments
Table 1: Comparison with stopping-specific baselines: TPCA, premature-stop rate, correctness rate, and overhead (GSM8K + HotpotQA, $n=500$ each, $\delta=0.1$ for EDFL).
Method TPCA Prem Stop Rate Correctness Overhead Formal $\delta$
GSM8K
   Answer-stability ($k=2$) 96.8 $\pm$ 3.1 0.142 $\pm$ 0.015 73.4% 2%
   Verifier-first 118.4 $\pm$ 4.2 0.089 $\pm$ 0.012 89.2% 28%
   Sequential-EDFL 84.3 $\pm$ 2.0 0.083 $\pm$ 0.010 76.8% 12%
HotpotQA
   Answer-stability ($k=2$) 108.2 $\pm$ 3.8 0.156 $\pm$ 0.018 69.8% 2%
   Verifier-first 132.7 $\pm$ 4.9 0.094 $\pm$ 0.013 85.6% 31%
   Sequential-EDFL 91.7 $\pm$ 2.7 0.091 $\pm$ 0.011 68.9% 12%
Answer-stability: stop when extracted answer unchanged for $k=2$ boundaries. Verifier-first: stop at first verifier pass. Premature-stop rate measured under sufficiency label (Section 4).
@@
[…]
Table 3: EDFL + Gate: Correctness improvement across benchmarks ($n=500$ each, $\delta=0.1$).
Setting Correctness HighLift+Incorrect TPCA Ovhd
EDFL (base) 76.8% 18.5% 90.7 12%
+ Sentence boundary only 79.6% 14.7% 93.2 13%
+ Boundary + verifier (recommended) 83.4% 10.9% 97.8 16%
We evaluate Sequential-EDFL on generation stopping across six benchmarks: GSM8K cobbe2021training (mathematical reasoning), HotpotQA yang2018hotpotqa (multi-hop QA), ASQA stelmakh2022asqa (long-form QA), TruthfulQA lin2022truthfulqa (factual accuracy), ProofWriter (logical reasoning), and LegalBench (legal analysis). Additional evaluation on code generation tasks (HumanEval, MBPP) appears in Appendix U.
Baselines: We compare against heuristic stopping methods: Fixed-length (150 tokens), Entropy threshold (stop when entropy $\tau_{H}$), Conformal stopping (calibration-set thresholds), SelfCheck manakul2023selfcheckgpt (self-consistency), and E-Value (fixed-$\lambda$ e-process). Additionally, we implement and evaluate two stopping-specific baselines: answer-stability stopping and verifier-first stopping (see Section 4.1).
Models: We evaluate on GPT-4 (gpt-4-0613), LLaMA-2-7B, and LLaMA-2-13B. Limitation: Our model suite focuses on models available at the time of experiments. For ACL’26, evaluation should ideally include at least one current open-source frontier model (e.g., Llama-3-70B, Mistral-Large, Qwen2.5) and one recent closed API model (e.g., GPT-4o, Claude 3.5 Sonnet, Gemini 1.5 Pro) to demonstrate transferability across architectures, training paradigms, and model families. This would validate that EDFL’s calibration properties (premature-stop rate tracking $\delta$) and efficiency gains (22-28% token reduction) generalize beyond the specific models evaluated here. Initial validation suggests the method is architecture-agnostic (it operates on probability distributions), but empirical confirmation across diverse models would strengthen the contribution. Verification mechanisms vary by domain: GSM8K uses symbolic calculator execution, HotpotQA and ASQA use search engine validation (top-5 results), TruthfulQA uses knowledge base fact-checking, ProofWriter uses theorem prover validation, and LegalBench uses legal database lookup.
For each benchmark, we randomly split examples 70% train, 10% calibration, 20% test. We train the correctness predictor on training data, compute conformal thresholds on calibration data, and report all metrics on held-out test sets. We set $\delta\in\{0.03,0.05,0.10\}$ to examine error-cost tradeoffs.
To evaluate $\delta$-control, we define an operational ground-truth label for sufficiency. For each sequence, we run generation to completion (until EOS or maximum length $T$), extracting the final answer $A_{T}$. At each candidate stop time $t$, we define the sequence as sufficient at time $t$ if the answer extracted at time $t$ matches $A_{T}$ (answer stability) and the sequence at time $t$ is syntactically complete (ends at sentence boundary or EOS). Type-I error (premature stop) occurs when we stop at time $\tau$ while the sequence is not sufficient at $\tau$. This operational definition aligns with the theoretical guarantee: we control the probability of stopping before information lift has accumulated sufficiently (as measured by answer stability). We validate this definition through LLM-as-judge agreement analysis and explore alternative sufficiency definitions (semantic stability, perplexity stability, information-theoretic) in Appendix S and Appendix T.
We report four key metrics. The empirical Type-I error (premature stop rate) measures the fraction of stopped sequences that violate the sufficiency label, serving as our primary calibration metric for $\delta$-control. The TPCA (Tokens Per Certified Answer) measures the mean number of generated tokens until a certification event (lower is better). The correctness rate measures the fraction of stopped sequences that produce factually correct answers, evaluated separately and not $\delta$-controlled. Finally, the correctness-sufficiency gap measures the percentage of high-lift sequences that are sufficient but incorrect.
We implement the standard e-value stopping rule using a fixed-$\lambda$ non-mixture e-process (no self-normalization), following the betting/e-process literature (vovk2021evalue; ramdas2023gametheoretic). We sweep $\lambda$ on a validation split and report test performance.
4.1 Comparison with Stopping-Specific Baselines
We implement and evaluate two natural stopping-specific baselines that target similar objectives through different mechanisms: answer-stability stopping and verifier-first stopping. These methods represent practical alternatives that directly target stopping criteria without requiring skeleton construction.
Answer-stability stopping extracts the answer string at each sentence boundary using task-specific parsing (numeric extraction for GSM8K, span extraction for QA, logical form extraction for ProofWriter). It stops when the extracted answer remains unchanged for $k=2$ consecutive sentence boundaries (tuned on validation data). Verifier-first stopping runs the lightweight verifier (arithmetic checker for GSM8K, retrieval overlap for HotpotQA/ASQA, etc.) continuously at each sentence boundary and stops when the verifier passes with confidence above a per-task calibrated threshold ($\tau_{c}=0.7$ as in Section 5).
Table 1 compares Sequential-EDFL against these stopping-specific baselines across TPCA, premature-stop rate (under our sufficiency label), correctness rate, and computational overhead. Sequential-EDFL achieves lower TPCA than answer-stability stopping (84.3 vs. 96.8 on GSM8K) while maintaining controlled premature-stop rate ($\delta=0.1$). Answer-stability stopping exhibits higher premature-stop rates (0.142 on GSM8K) because it may stop when models oscillate between incorrect but stable answers, failing to distinguish sufficiency from incorrect stability. Verifier-first stopping achieves higher correctness rates (89.2% vs. 76.8% on GSM8K) but at the cost of substantially higher TPCA (118.4 vs. 84.3) and overhead (28% vs. 12%) due to continuous verifier calls. Sequential-EDFL’s unique advantage is the formal $\delta$-control for sufficiency violations (Theorem 3.1), which neither baseline provides, combined with competitive efficiency (lower TPCA than both baselines) and moderate overhead.
Sequential-EDFL achieves lower TPCA than both stopping-specific baselines while maintaining controlled premature-stop rate at $\delta=0.1$. Answer-stability stopping’s higher premature-stop rates (0.142 on GSM8K, 0.156 on HotpotQA) reflect its inability to distinguish sufficient from insufficient stopping when models stabilize on incorrect answers. Verifier-first stopping prioritizes correctness but requires substantially more tokens (118.4 vs. 84.3 on GSM8K) and overhead (28% vs. 12%) due to continuous verifier evaluation. EDFL’s formal $\delta$-control provides a unique guarantee that neither baseline offers, while achieving superior efficiency.